OpenStudy (anonymous):

Could someone please explain how is sqrt(40)^3 = ( 2^3 * sqrt(10)^3 ) ?

OpenStudy (anonymous):

40 = 4 * 10 SQRT(40)^3 = SQRT(4*10)^3 = SQRT(4)^3 * SQRT(10)^3 = 2^3 * SQRT(10)^3